OpenStudy (anonymous):

what is the solution set of the equation 2x^2 - 6x - 8 = 0

OpenStudy (yolomcswagginsggg):

Let's solve your equation step-by-step. 2x2−6x−8=0 Step 1: Factor left side of equation. 2(x+1)(x−4)=0 Step 2: Set factors equal to 0. x+1=0 or x−4=0 x=−1 or x=4 Answer: x=−1 or x=4

OpenStudy (yolomcswagginsggg):

I hope that helped!

OpenStudy (anonymous):

so the answer would be -1,-4? or just one or the other?

OpenStudy (yolomcswagginsggg):

Just one depending on what the answers are, unless its multiple choice ;p

OpenStudy (anonymous):

I have to fill it in :( which one would you go for?

OpenStudy (yolomcswagginsggg):

hmm....

OpenStudy (yolomcswagginsggg):

i'm not sure....

OpenStudy (yolomcswagginsggg):

@welshfella what would you choose?

OpenStudy (anonymous):

it also has, The solution set is {-1,[1]} sorry

OpenStudy (yolomcswagginsggg):

Oh XD

OpenStudy (yolomcswagginsggg):

Ok, then XDDD

OpenStudy (yolomcswagginsggg):

I fanned you ;p

OpenStudy (anonymous):

thanks :) would it be -1 then

OpenStudy (yolomcswagginsggg):

Alrighty just do @yolomcswagginsgg when you need halp ;p